# Reward-Forcing: Autoregressive Video Generation with Reward Feedback

### v1

@@ 1 Introduction
Diffusion models [18, 33] have emerged as a powerful class of generative models, achieving state-of-the-art results in a wide range of domains, including image synthesis, audio generation, and molecular modeling. By iteratively denoising data from a simple noise distribution, these models are capable of producing high-fidelity samples that capture complex data distributions. Their theoretical foundation rooted in stochastic differential equations, combined with their empirical robustness, has made diffusion models a dominant approach in the generative modeling landscape.
Building on this success, video diffusion models [19, 3] extend the capabilities of diffusion-based generation to the temporal domain. Unlike images, videos require the model to generate spatially coherent and temporally consistent sequences of frames, which significantly increases the modeling complexity. Recent advancements have adapted diffusion models to handle temporal correlations by incorporating spatiotemporal attention, motion priors, and multi-frame conditioning mechanisms. These models have demonstrated promising results in generating short to medium-length video clips with high perceptual quality.
However, most existing video diffusion models are built upon bidirectional architectures [55, 29], where the generation of each frame depends on information from both past and future timesteps. While effective for high quality video generation, bidirectional models are inherently unsuitable for streaming or real-time video generation, taking several minutes to generate a 5-second video. Recent efforts have explored transforming bidirectional video diffusion models into autoregressive formulations [60, 26] to enable low-latency and scalable generation. These methods typically rely on distillation from powerful bidirectional teachers, but their performance is often limited by the quality of the teacher model and the challenges of maintaining temporal coherence in the absence of future context.
Meanwhile, reinforcement learning (RL) has emerged as a powerful paradigm for post-training fine-tuning for enhancing the quality and alignment of diffusion-based image generation. By optimizing generation policies directly with respect to reward signals—such as human preferences, aesthetic scores, or perceptual quality—RL enables fine-grained control over generative behavior beyond supervised objectives. Techniques such as Reinforcement Learning with Human Feedback (RLHF) [40, 32] have demonstrated success in aligning language and vision-language models, and recent works have begun applying RL to tune diffusion models for improved fidelity, diversity, or alignment with desired styles.
@@
  • We demonstrate that the performance of existing methods for converting bidirectional video diffusion models into autoregressive models are bounded by the teacher model’s performance.
  • We observe that consistent motions are learned before texture, consistent with previous studies on image generation, and propose a framework that uses pure reward signals to guide autoregressive generation of high-quality videos.
  • Extensive experiments on VBench show that our method achieves video quality comparable to baseline autoregressive models while maintaining competitive overall scores.
@@ 2 Related work
Denoising Diffusion Probabilistic Models (DDPM) pioneered modern image generation by casting generative modeling as a Markovian noising–denoising process whose reverse dynamics are learned with a simple, noise‑conditional score network [18]. While DDPMs achieve impressive sample quality, they rely on hundreds or thousands of sequential denoising steps. Denoising Diffusion Implicit Models (DDIM) alleviate this inefficiency by interpreting the learned stochastic process as a non‑Markovian deterministic ordinary differential equation, enabling much faster inference without retraining and preserving visual fidelity [50]. Subsequent flow‑matching methods such as the Flow Matching framework further unify diffusion and continuous‑normalizing‑flow viewpoints by directly training vector fields that map noise to data in a single pass, often reducing both training variance and sampling cost [30, 33]. The most widely used architecture for such tasks was UNet [44], which was later revolutionized by the introduction of DiT [41]. Due to the high computational cost of operating directly in pixel space, modern architectures typically first project the input into a latent space [43, 1], where subsequent computations are performed. The final results are then decoded back into pixel space using a decoder [27].
@@
Video diffusion models represent a significant advancement in generative AI, extending the principles of image diffusion models to handle temporal dynamics for tasks such as text-to-video generation, image-to-video synthesis, and video editing. Early extensions to video, such as Video Diffusion Models [19], introduced 3D U-Net architectures with factorized spatio-temporal attention to ensure temporal coherence in video output with relative position embeddings [46]. Subsequent models like Imagen Video [17] employed cascaded diffusion pipelines for high-definition videos, while Make-A-Video [49] leveraged pretrained text-to-image models and unsupervised video data to learn visual appearance and motion without requiring paired text-video data. By introducing novel spatial-temporal modules and a multi-stage video generation pipeline, these methods achieve state-of-the-art results in video quality, temporal consistency, and alignment with textual input. Latent space approaches, exemplified by models such as Stable Video Diffusion [1, 2], enhanced efficiency by operating on compressed representations and incorporating temporal attention mechanisms. More recent innovations include zero-shot methods like Text2Video-Zero [25], which enable video generation without video-specific training, and Sora [39], utilizing Diffusion Transformers on spacetime patches for scalable, high-fidelity outputs. These developments address challenges in temporal consistency and computational demands, paving the way for applications in entertainment, simulation, and content creation. The quality of generation is further enhanced by following models such as Wan [54], HunyuanVideo [29], and Veo [15].
@@
Although diffusion models are able to perform high quality image and video synthesis, they suffer from extremely high denoising steps with high cost. In order to solve the problem, one direction is to design better ODE solvers which enable few-step sampling such as DPM-Solver [35], GENIE [12], and S4S [14]. An alternative approach involves distillation techniques to compress multi-step diffusion processes into fewer or single-step generations, including Score Distillation Sampling (SDS) for leveraging pre-trained diffusion priors in optimization tasks [36], Distribution Matching Distillation [59] which transforms diffusion models into efficient one-step generators with minimal quality loss, its enhanced successor DMD2 [58] that further improves training efficiency and performance, and Consistency Models that enable high-quality one- or few-step sampling by learning consistent mappings from noise to data [52].
@@
While most diffusion models are trained with full-sequence denoising, they usually require substantially long time to generate videos up to several seconds. Thus autoregressive and streaming variants have been proposed to enable fast frame-by-frame or chunk-by-chunk generation  [60, 20]. Among these works, FIFO-Diffusion [26] proposes a training-free method that uses a queue to hold frames of different noise levels with the frames closer to the beginning of the queue have lower noise levels and require fewer denoising steps, whereas frames at the end of the queue have higher noise levels and require more. The method is able to generate substantially longer videos than the original model. However, due to the heterogeneous denoising steps in the input, the generated videos often show degraded quality and inconsistency. Ouroboros-Diffusion [7] tries to solve the inconsistency problem by introducing a novel latent sampling technique at the end of the queue with subject-aware cross-frame attention mechanism. The heterogeneous noise level used by [26, 7] shares the same idea as Diffusion-Forcing [5, 51] where the models are trained to perform denoising on noisy frames with different noise levels. However, due to the large combination of noise levels, recent models [6] usually first train the model with same noise level, then finetune it to handle different noise levels. Later works [28] try to distill the model into few-step generators which alleviates the problem. Another line of works such as CausVid [60] and Self Forcing [20] adopts the same autoregressive manner but with same noise levels which show improved video quality.
@@
Reinforcement learning has been widely applied in Large Language Models [39, 13, 10] to finetune models to align better with user preferences [9, 42, 31]. Such techniques have also been generalized to diffusion models. ImageReward [57] builds the first general-purpose text-to-image human preference reward model which can be used in the Reward Feedback Learning (ReFL) framework to effectively align image generation with human preferences. VisionReward [56] extends it to both images and videos by designing a fine-grained, multi-dimensional reward model, achieving new state-of-the-art performances on various benchmark datasets. VideoAlign [32] introduces a VLM-based reward model to address three critical dimensions including Visual Quality, Motion Quality and Text Alignment. ROCM [47] proposes a direct reward optimization framework for applying reinforcement learning from human feedback (RLHF) to consistency models, enabling efficient training without the need for policy gradients. Reward-Instruct [37] achieves fast image synthesis through reward-centric approaches.
Motion in Video Diffusion Models
Motion modeling remains a central challenge in video diffusion. Many early video diffusion models, such as Video Diffusion Models [19] primarily focus on spatiotemporal attention mechanisms but treat motion implicitly. More recent efforts like MCDiff [8] and VideoJAM [4] attempt to model motion more explicitly, using a flow completion model or joint appearance-motion representations. And Motion-I2V [48] further explores motion disentanglement in the image-to-video task, suggesting a broader trend toward treating motion and appearance as separable generative components.
@@ 3.1 Preliminaries
$q\!\left(x_{t}\mid x_{0}\right)$ $=\mathcal{N}\!\bigl(\alpha_{t}x_{0},\;\sigma_{t}^{2}\mathbf{I}\bigr),$ (1)
@@
$\mathrm{d}x_{t}=\bigl[f(x_{t},t)-g(t)^{2}\,\nabla_{x_{t}}\log q_{t}(x_{t})\bigr]\mathrm{d}t+g(t)\,\mathrm{d}\bar{w}_{t}.$ (2)
@@
While DDPMs rely on stochastic trajectories, flow matching [30] learns a deterministic velocity field that continuously transports a tractable prior $p_{0}$ to the data distribution $p_{1}$. Given a coupling $(x_{0},x_{1})\sim p_{0}\times p_{1}$, define linear interpolates $x_{t}=(1-t)\,x_{0}+t\,x_{1}$. The ground-truth velocity $v^{\star}(x_{t},t)$ is
$v^{\star}(x_{t},t)=\frac{dx_{t}}{dt}=x_{1}-x_{0}.$ (3)
@@
$\mathcal{L}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1)}\,\mathbb{E}_{(x_{0},x_{1})}\Bigl[\,\bigl\|v_{\theta}(x_{t},t)-v^{\star}(x_{t},t)\bigr\|^{2}\Bigr].$ (4)
At inference, samples evolve deterministically via the ODE $\dot{x}_{t}=v_{\theta}(x_{t},t)$, often requiring far fewer steps than stochastic diffusion samplers while maintaining high generative fidelity. In this work, we use Wan2.1 [55] which is trained using the flow matching objective.
@@
Same as CausVid [60] and SelfForcing [20], we also utilize the distilled few-step model as the starting model for faster generation. It is achieved by minimizing the reverse KL divergence between data distribution and stduent generator’s output distribution which can be formulated as
$\nabla_{\theta}\mathcal{L}_{\mathrm{DMD}}$ $=\mathbb{E}_{t}\!\bigl[\nabla_{\theta}\operatorname{KL}\bigl(p_{\mathrm{fake},t}\|p_{\mathrm{real},t}\bigr)\bigr]$ (5)
@@ 3.2 Training with ODE and Reward Guidance
Recent studies have shown that diffusion models typically generate a global coarse structure before progressively refining texture details [8, 23, 53, 34]. We empirically observe a similar phenomenon when converting bidirectional video diffusion teachers into autoregressive models which can be seen in Fig. 4 here. After distilling the model into few-step models, similar to CausVid [60], we use sample ODE trajectories for early training. The process is done by first sample noise inputs $\{x_{T}^{i}\}_{i=1}^{L}$ from $\mathcal{N}(0,I)$ and then use an ODE solver with the pretrained teacher to generate reverse trajectories $\{x_{t}^{i}\}_{i=1}^{L}$ across all timesteps. The model is then trained with the following objective:
$\mathcal{L}_{\text{ode}}=\mathbb{E}_{x,t^{i}}\left\|G_{\phi}(\{x_{t^{i}}^{i}\},\{t^{i}\})-\{x_{0}^{i}\}\right\|^{2}.$ (6)
@@
We also observe that ODE initialization provides a strong motion prior, allowing reward supervision to refine texture without destabilizing temporal dynamics.

### v2

@@ 1 Introduction
Diffusion models (Ho et al., 2020; Liu et al., 2023) have emerged as a powerful class of generative models, achieving state-of-the-art results in a wide range of domains, including image synthesis, audio generation, and molecular modeling. By iteratively denoising data from a simple noise distribution, these models are capable of producing high-fidelity samples that capture complex data distributions. Their theoretical foundation rooted in stochastic differential equations, combined with their empirical robustness, has made diffusion models a dominant approach in the generative modeling landscape.
Building on this success, video diffusion models (Ho et al., 2022b; Brooks et al., 2024) extend the capabilities of diffusion-based generation to the temporal domain. Unlike images, videos require the model to generate spatially coherent and temporally consistent sequences of frames, which significantly increases the modeling complexity. Recent advancements have adapted diffusion models to handle temporal correlations by incorporating spatiotemporal attention, motion priors, and multi-frame conditioning mechanisms. These models have demonstrated promising results in generating short to medium-length video clips with high perceptual quality.
However, most existing video diffusion models are built upon bidirectional architectures (Wang et al., 2025; Kong et al., 2024), where the generation of each frame depends on information from both past and future timesteps. While effective for high quality video generation, bidirectional models are inherently unsuitable for streaming or real-time video generation, taking minutes to generate a single video with 5 seconds. Recent efforts have explored transforming bidirectional video diffusion models into autoregressive formulations (Yin et al., 2025; Kim et al., 2024) to enable low-latency and scalable generation. These methods typically rely on distillation from powerful bidirectional teachers, but their performance is often limited by the quality of the teacher model and the challenges of maintaining temporal coherence in the absence of future context.
In parallel, reinforcement learning (RL) has gained traction as a post-training strategy for enhancing the quality and alignment of diffusion-based image generation. By optimizing generation policies directly with respect to reward signals—such as human preferences, aesthetic scores, or perceptual quality—RL enables fine-grained control over generative behavior beyond supervised objectives. Techniques such as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022; Liu et al., 2025a) have demonstrated success in aligning language and vision-language models, and recent works have begun applying RL to tune diffusion models for improved fidelity, diversity, or alignment with desired styles.
@@
  • We show that the performance of existing methods for converting bidirectional video diffusion models into autoregressive models are bounded by the teacher’s performance.
  • We observe that consistent motions are learned before texture which aligns with previous studies on image generation and propose a framework that uses pure reward signals to guide autoregressive generation of high quality videos.
  • Extensive experiments on VBench show that our proposed method is able to generate high quality videos, comparable to baseline autoregressive video diffusion models on quality and total score.
@@ 2 Related work
Denoising Diffusion Probabilistic Models (DDPM) pioneered the modern image generation by casting generative modeling as a Markovian noising–denoising process whose reverse dynamics are learned with a simple, noise‑conditional score network (Ho et al., 2020). While DDPMs achieve impressive sample quality, they rely on hundreds or thousands of sequential denoising steps. Denoising Diffusion Implicit Models (DDIM) alleviate this inefficiency by interpreting the learned stochastic process as a non‑Markovian deterministic ordinary differential equation, enabling much faster inference without retraining and preserving visual fidelity (Song et al., 2021). Subsequent flow‑matching methods such as the Flow Matching framework further unify diffusion and continuous‑normalizing‑flow viewpoints by directly training vector fields that map noise to data in a single pass, often reducing both training variance and sampling cost (Lipman et al., 2022; Liu et al., 2023). The most widely used architecture for such tasks was UNet (Ronneberger et al., 2015), which was later revolutionized by the introduction of DiT (Peebles and Xie, 2023). Due to the high computational cost of operating directly in pixel space, modern architectures typically first project the input into a latent space (Rombach et al., 2022; Blattmann et al., 2023a), where subsequent computations are performed. The final results are then decoded back into pixel space using a decoder (Kingma and Welling, 2014).
@@
Video diffusion models represent a significant advancement in generative AI, extending the principles of image diffusion models to handle temporal dynamics for tasks such as text-to-video generation, image-to-video synthesis, and video editing. Early extensions to video, such as Video Diffusion Models (Ho et al., 2022b), introduced 3D U-Net architectures with factorized spatio-temporal attention to ensure temporal coherence in video output with relative position embeddings (Shaw et al., 2018). Subsequent models like Imagen Video (Ho et al., 2022a) employed cascaded diffusion pipelines for high-definition videos, while Make-A-Video (Singer et al., 2023) leveraged pretrained text-to-image models and unsupervised video data to learn visual appearance and motion without requiring paired text-video data. By introducing novel spatial-temporal modules and a multi-stage video generation pipeline, it achieves state-of-the-art results in video quality, temporal consistency, and alignment with textual input. Latent space approaches, exemplified by models such as Stable Video Diffusion (Blattmann et al., 2023a; b), enhanced efficiency by operating on compressed representations and incorporating temporal attention mechanisms. More recent innovations include zero-shot methods like Text2Video-Zero (Khachatryan et al., 2023), which enable video generation without video-specific training, and Sora (OpenAI, 2023), utilizing Diffusion Transformers on spacetime patches for scalable, high-fidelity outputs. These developments address challenges in temporal consistency and computational demands, paving the way for applications in entertainment, simulation, and content creation. The quality of generation is further enhanced by following models such as Wan (Wan et al., 2025), HunyuanVideo (Kong et al., 2024) and Veo (Google, 2025).
@@
Although diffusion models are able to perform high quality image and video synthesis, they suffer from extremely high denoising steps with high cost. In order to solve the problem, one direction is to design better ODE solvers which enable few-step sampling such as DPM-Solver (Lu et al., 2022), GENIE (Dockhorn et al., 2022) and S4S (Frankel et al., 2025). An alternative approach involves distillation techniques to compress multi-step diffusion processes into fewer or single-step generations, including Score Distillation Sampling (SDS) for leveraging pre-trained diffusion priors in optimization tasks (Luo et al., 2024), Distribution Matching Distillation (Yin et al., 2024b) which transforms diffusion models into efficient one-step generators with minimal quality loss, its enhanced successor DMD2 (Yin et al., 2024a) that further improves training efficiency and performance, and Consistency Models that enable high-quality one- or few-step sampling by learning consistent mappings from noise to data (Song et al., 2023).
@@
While most diffusion models are trained with full-sequence denoising, they usually require substantially long time to generate videos up to several seconds. Thus autoregressive and streaming variants have been proposed to enable fast frame-by-frame or chunk by chunk generation  (Yin et al., 2025; Huang et al., 2025). Among the work, FIFO-Diffusion (Kim et al., 2024) proposes a training-free method that uses a queue to hold frames of different noise levels with the frames closer to the beginning of the queue having lower noise level which requires less denoising and frames closer to the end of the queue have high noise levels which requires more denoising steps. The method is able to generate substantially longer videos than the original model. However, due to the heterogeneous denoising steps in the input, the generated videos often show degraded quality and inconsistency. Ouroboros-Diffusion (Chen et al., 2025b) tries to solve the inconsistency problem by introducing a novel latent sampling technique at the end of the queue with subject-aware cross-frame attention mechanism. The heterogeneous noise level used by Kim et al. (2024); Chen et al. (2025b) shares the same idea as Diffusion-Forcing (Chen et al., 2024; Song et al., 2025) where the models are trained to perform denoising on noisy frames with different noise levels. However, due to the large combination of noise levels, recent models (Chen et al., 2025a) usually first train the model with same noise level, then finetune it to handle different noise levels. Later work (Kodaira et al., 2025) tries to distill the model into few-step generators which alleviates the problem. Another line of works such as CausVid (Yin et al., 2025) and Self Forcing (Huang et al., 2025) adopts the same autoregressive manner but with same noise levels which show improved video quality.
@@
Reinforcement learning has been widely applied in Large Language Models OpenAI (2023); Dubey et al. (2024); Comanici et al. (2025) to finetune models to align better with user preferences Christiano et al. (2017); Rafailov et al. (2023); Liu et al. (2024). Such techniques have also been generalized to diffusion models. ImageReward Xu et al. (2023) builds the first general-purpose text-to-image human preference reward model which can be used in the Reward Feedback Learning (ReFL) framework to effectively align image generation with human preferences. VisionReward Xu et al. (2024) extends it to both images and videos by designing a fine-grained, multi-dimensional reward model, achieving new state-of-the-art performances on various benchmark datasets. VideoAlign Liu et al. (2025a) introduces a VLM-based reward model to address three critical dimensions including Visual Quality, Motion Quality and Text Alignment. ROCM Shekhar and Zhang (2025) proposes a direct reward optimization framework for applying reinforcement learning from human feedback (RLHF) to consistency models, enabling efficient training without the need for policy gradients. Reward-Instruct Luo et al. (2025) achieves fast image synthesis through reward-centric approaches.
@@ 3.1 Preliminaries
$q\!\left(x_{t}\mid x_{0}\right)$ $=\mathcal{N}\!\bigl(\alpha_{t}x_{0},\;\sigma_{t}^{2}\mathbf{I}\bigr),$
@@
$\mathrm{d}x_{t}=\bigl[f(x_{t},t)-g(t)^{2}\,\nabla_{x_{t}}\log q_{t}(x_{t})\bigr]\mathrm{d}t+g(t)\,\mathrm{d}\bar{w}_{t}.$
@@
While DDPMs rely on stochastic trajectories, flow matching (Lipman et al., 2022) learns a deterministic velocity field that continuously transports a tractable prior $p_{0}$ to the data distribution $p_{1}$. Given a coupling $(x_{0},x_{1})\sim p_{0}\times p_{1}$, define linear interpolates $x_{t}=(1-t)\,x_{0}+t\,x_{1}$. The ground-truth velocity $v^{\star}(x_{t},t)$ is
$v^{\star}(x_{t},t)=\frac{dx_{t}}{dt}=x_{1}-x_{0}.$
@@
$\mathcal{L}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1)}\,\mathbb{E}_{(x_{0},x_{1})}\Bigl[\,\bigl\|v_{\theta}(x_{t},t)-v^{\star}(x_{t},t)\bigr\|^{2}\Bigr].$
At inference, samples evolve deterministically via the ODE $\dot{x}_{t}=v_{\theta}(x_{t},t)$, often requiring far fewer steps than stochastic diffusion samplers while maintaining high generative fidelity. In this work, we use Wan2.1 (Wang et al., 2025) which is trained using the flow matching objective.
@@
Same as CausVid (Yin et al., 2025) and SelfForcing (Huang et al., 2025), we also utilize the distilled few-step model as the starting model for faster generation. It is achieved by minimizing the reverse KL divergence between data distribution and stduent generator’s output distribution which can be formulated as
$\nabla_{\theta}\mathcal{L}_{\mathrm{DMD}}$ $=\mathbb{E}_{t}\!\bigl[\nabla_{\theta}\operatorname{KL}\bigl(p_{\mathrm{fake},t}\|p_{\mathrm{real},t}\bigr)\bigr]$
@@ 3.2 Training with ODE and Reward Guidance
Recent studies have shown that diffusion models typically generate a global coarse structure before progressively refining texture details (Chen et al., 2023; Jiang et al., 2024; Sun et al., 2025; Liu et al., 2025b). We empirically observe a similar phenomenon when converting bidirectional video diffusion teachers into autoregressive models which can be seen in Fig. 3 here. After distilling the model into few-step models, similar to CausVid (Yin et al., 2025), we use sample ODE trajectories for early training. The process is done by first sample noise inputs $\{x_{T}^{i}\}_{i=1}^{L}$ from $\mathcal{N}(0,I)$ and then use an ODE solver with the pretrained teacher to generate reverse trajectories $\{x_{t}^{i}\}_{i=1}^{L}$ across all timesteps. The model is then trained with the following objective:
$\mathcal{L}_{\text{ode}}=\mathbb{E}_{x,t^{i}}\left\|G_{\phi}(\{x_{t^{i}}^{i}\},\{t^{i}\})-\{x_{0}^{i}\}\right\|^{2}.$
@@
Optimization with Reward Feedback
After initializing the model with an ODE-based process that teaches motion synthesis, we introduce a reward-guided optimization stage to enhance video quality. Specifically, we adopt ImageReward (Xu et al., 2023) as our reward model and incorporate it into training in a differentiable manner to directly guide the video diffusion model.
Let the generated video be denoted as
$\hat{x}_{1:T}=G_{\theta}(z)$
, where $G_{\theta}$ is the autoregressive video generator parameterized by $\theta$, and $z$ is the latent input. The reward model $\mathcal{R}(\cdot)$ assigns a scalar reward indicating perceptual quality. We define the reward-guided objective as:
$\mathcal{L}_{\text{reward}}(\theta)=-\mathbb{E}_{z\sim\mathcal{Z}}\left[\mathcal{R}(\hat{x}_{T})\right].$
3.3 Baseline Methods
We benchmark three classes of approaches: (i) standard bidirectional diffusion models, (ii) autoregressive methods that render videos one (latent) frame at a time, and (iii) autoregressive methods that synthesize videos in temporal chunks. For standard diffusion models, we include LTX-Video (HaCohen et al., 2024) and Wan2.1 (Wan et al., 2025) which is the base model used by Yin et al. (2025); Huang et al. (2025) and our method. For frame-wise autoregressive models, we include NOVA (Deng et al., 2025) which formulates the video generation problem as non-quantized autoregressive modeling of temporal frame-by-frame prediction and spatial set-by-set prediction, Pyramid Flow (Jin et al., 2025) that interprets the original denoising trajectory as a series of pyramid stages and Self Forcing (Huang et al., 2025) that trains the autoregressive models by relying on self-generated frames instead of ground-truth ones. For Chunk-wise models, we include SkyReels-V2 (Chen et al., 2025a), MAGI-1(Sand-AI, 2025) that utilizes transformer based VAE and increased noise levels among chunks that are being generated, CausVid (Yin et al., 2025) that uses causal attention at training time and self-rollout at inference time and Self Forcing (Huang et al., 2025) which utilizes similar architectures as its frame-wise counterpart.